OpenStudy (anonymous):

Oneee more lol Find the coordinates of the midpoint of the segment whose endpoints are (10, 6) and (-4, 8). @Yttrium

OpenStudy (anonymous):

I'm horrible at remembering the formulas because I just started this advanced school and didnt learn any of this before so I'm just kinda stuck! Lol.

OpenStudy (yttrium):

It's easy. Just get the averages of coordinates at x and y and that's it. :)

hartnn (hartnn):

at a time, practice few problems which use same formula! you'll remember that formula for long time :)

OpenStudy (yttrium):

In short, add them and just divide by two. Separate from x and y and you will form the coordinates of it's midpoint.

hartnn (hartnn):

yeah, an if you want the equation of the formula, The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

So basically \[\frac{ x1 + x2 }{ 2 } and \frac{ x1+y1 }{ 2 }\]

OpenStudy (anonymous):

?

OpenStudy (yttrium):

Yup @CrayolaCrayon_

OpenStudy (anonymous):

Ahh I thought so!

hartnn (hartnn):

(y1+y2)/2 *** but i think you meant that only

OpenStudy (anonymous):

Oh yes, that is what I meant!

hartnn (hartnn):

so what mid-point you got for (10, 6) and (-4, 8). ??

OpenStudy (anonymous):

so it's \[\frac{ 10+6 }{ 2 } and \frac{ -4+8 }{ 2 }\] ?

hartnn (hartnn):

nopes, what u did is you added x and y co-ordinates you needed to add x co-ordinates of both points

hartnn (hartnn):

x1 = 10 x2 = -4 x= ... ?

OpenStudy (anonymous):

Oh oops wait!

OpenStudy (anonymous):

\[\frac{ 10-4}{ 2} and \frac{ 6+8 }{ }\]

OpenStudy (anonymous):

oops, forgot the 2 lol

hartnn (hartnn):

yes, and those = ... ? simplify

OpenStudy (anonymous):

\[\frac{ 16 }{ 2 } and \frac{ 14 }{ 2 } \rightarrow \frac{ 8 }{ 1 } and \frac{ 7 }{ 1 } \rightarrow\] 8 and 7

hartnn (hartnn):

where does 16 come from ? 10-4 = ... ?

OpenStudy (anonymous):

oopsss! I'm so tired I can't think straight lol. I accidentally added. Make that 3 and 7 haha

OpenStudy (anonymous):

(3,7)

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^